OpenStudy (anonymous):

A 13-foot ladder is placed so that it reaches to a point on the wall that is 2 feet higher than twice the distance from the base of the wall to the base of the ladder. (a) How far from the wall is the base of the ladder? In Feet. (b) How high does the ladder reach? In Feet.

OpenStudy (anonymous):

i gotta make an equation like c^2=a^2 + b^2, right?

OpenStudy (anonymous):

i much rather have you not tell me the answer :P

OpenStudy (anonymous):

so its 13^2= 2+2x^2 + x^2?

myininaya (myininaya):